OpenStudy (anonymous):

How do you solve this system? 2x-3y=1 x+y=3

OpenStudy (anonymous):

u can solve by substitution or elimination... by sub let y=-x+3 then 2x-3(-x+3)=1

OpenStudy (anonymous):

in this case it wants me to solve it graphically?

OpenStudy (anonymous):

then set both equations in slope intercept form and find the intersection

hero (hero):

Find the x and y intercepts of each line

OpenStudy (anonymous):

how do you find the intercepts?

OpenStudy (anonymous):

so y=-x+3 and y=2/3x+1/3

OpenStudy (anonymous):

just graph both lines and the solution is the place where they intersect

OpenStudy (anonymous):

oh ok thank you!!!

OpenStudy (anonymous):

so should be (8/5, 7/5)